OpenStudy (anonymous):

Need Help!! What is the solution to the system of equations? 2x+y=-3 x-2y=-4

OpenStudy (anonymous):

there are to different equations

OpenStudy (anonymous):

@campbell_st @CausticSyndicalist @cwrw238 @StudyGurl14

OpenStudy (anonymous):

i need it ASAP

OpenStudy (studygurl14):

2x + y = -3 x - 2y = -4 --> x = 2y - 4 Plug into the other equation 2(2y - 4) + y = -3 Distribute 4y - 8 + y= -3 Can you do the rest from here @MAli13chineta ?

OpenStudy (anonymous):

umm..... probably not

OpenStudy (studygurl14):

4y - 8 + y = -3 5y - 8 = -3 Solve for y @MAli13chineta

OpenStudy (anonymous):

@campbell_st what would the steps be to solve for y

OpenStudy (anonymous):

@StudyGurl14